OpenStudy (anonymous):

How many inches are in a 7/8 of a yard?

OpenStudy (campbell_st):

there are 36 inches in a yard. so its 7/8 times 36

OpenStudy (anonymous):

Thanks!

OpenStudy (unklerhaukus):

\[\frac 78 [\text {yd}]\]\[=\frac 78 [\text {yd}]\times\frac{3[\text {ft}]}{[\text {yd}]}\]\[=\frac 78 [\text {yd}]\times\frac{3[\text {ft}]}{[\text {yd}]}\times\frac{12[\text {in}]}{[\text {ft}]}\]\[=\frac 78\times3\times12[\text{in}]\]

OpenStudy (anonymous):

Thanks! This really helped!